OpenStudy (anonymous):

How do I evaluate csc(-7pie/6)

OpenStudy (anonymous):

\[\csc(\theta)=\frac{1}{\sin(\theta)}\]

OpenStudy (anonymous):

what about cot(pi)

OpenStudy (anonymous):

\[\cot(\theta)=\frac{1}{\tan(\theta)}\]

hero (hero):

Let \(\pi = 180^{\circ}\), then find the corresponding angle

OpenStudy (anonymous):

so would tan be undefined since tan(pi) has no trig function

OpenStudy (anonymous):

tan(pi)=0, cotan(pi)= undefined

OpenStudy (anonymous):

i have 1/tan(pi) , is this undefined?

OpenStudy (anonymous):

cos(pi)=1 and sin(pi)=0, and maths breaks when you try to divide by 0

OpenStudy (anonymous):

Yes, as tan(pi)=0, and dividing by 0 doesn't work.

OpenStudy (anonymous):

awesome so how would sec(3pi/4) work?

OpenStudy (anonymous):

\[\sec(\theta)=\frac{1}{\cos(\theta)}\]

OpenStudy (anonymous):

okay so after i have 1/cos(3(pi)/4) what do i do? right now i have 1/-\[-\sqrt{2}\]/2

OpenStudy (anonymous):

i have \[1/-\sqrt{2}/2\]

OpenStudy (anonymous):

if sinx=-2/5 and x is in the third quadrant, what are the other five trig functions of x